# Appearance-based Refinement for Object-Centric Motion Segmentation

### v1

@@ 3.4. Sim2Real training procedure
3.4.1 Pre-training on synthetic videos
@@
$L_{\text{corrector}}=\lambda_{\text{mask}}L_{\text{mask}}+\lambda_{\text{feat-recon}}L_{\text{feat-recon}}\vspace{-0.05cm}$ (8)
@@
3.4.2 Self-supervised adaptation to real-world videos
The Sim2Real domain gap is mitigated by self-supervised training on target real-world datasets. Specifically, for each real-world video dataset, we first employ our synthetically-trained selection-correction model to generate segmentation masks starting from the OCLR-flow proposals. The resultant masks are then treated as pseudo-labeling to fine-tune the mask corrector and the OCLR-flow segmentation model.
Adapting the mask correction model. The overall loss for fine-tuning remains the same as for synthetic training (Eq. 8). However, for adaptation, the refined masks generated by the synthetically-trained corrector (i.e., {~$\{\tilde{\mathcal{M}}_{i}^{k}\}$) are treated as pseudo-GT, and become the target masks in the loss. Note that, we only apply pseudo-labeling at the frames where exemplar masks present due to their high quality. Apart from the self-supervised training at the dataset level, we also employ a test-time per-sequence adaptation to further boost the performance. A detailed description is given in Section A.3.
@@ 3.4.2 Self-supervised adaptation to real-world videos
4. Experiments
4.1. Datasets and metrics
@@
Multi-object benchmark. We adopt the DAVIS2017-motion dataset (short form “DAVIS17-m”) that is proposed in [92] with jointly moving objects re-annotated. We further formulate a YouTubeVOS2018-motion dataset (short form “YTVOS18-m”), which is a subset of YTVOS18 [93], containing $120$ sequences that involve up to three objects. Additional information regarding this new benchmark can be found in Section B.3.
@@
Evaluation metrics. We benchmark our single object video segmentation performance using region similarity ($\mathcal{J}$), aka. intersection-over-union. For multi-object video segmentation, we additionally report the contour similarity ($\mathcal{F}$[55].
4.2. Implementation details
For pre-processing, we apply RAFT [79] to estimate optical flow with frame gaps ±$\pm 1$ (except for FBMS and YTVOS18-m with frame gaps ±$\pm 3$ owing to relatively slow video motions). All optical flows and RGB frame inputs are resized to $128\times 224$. The motion segmentation model proposes $K=3$ objects for each sequence, from which $t=10$ exemplar masks are selected for each object. Also, we use self-supervised pre-trained DINO transformers [9], ViT-S/16 for mask selector, ViT-S/8 for corrector, resulting in DINO feature resolutions $8\times 14$ and $16\times 28$, respectively. As a final post-processing step, we apply Conditional Random Field (CRF) segmentation to further refine the output masks. We refer the reader to Appendix A for more details regarding the model architecture, training settings, and inference time analysis.
@@ 4.2. Implementation details
${}^{{\dagger}}$ideoCutLER [88] un-sup. $\checkmark$ $480\times 854$ $57.0$ $59.0$ $55.1$ $57.3$ $57.4$ $57.2$
@@
${}^{{\dagger}}$orrFlow [40] semi-sup. $\checkmark$ $480\times 854$ $55.2$ $60.0$ $50.4$ $54.0$ $54.2$ $53.7$
${}^{{\dagger}}$ [46] semi-sup. $\checkmark$ $480\times 854$ $66.9$ $70.4$ $63.4$ $65.5$ $66.2$ $64.7$
${}^{{\dagger}}$ [30] semi-sup. $\checkmark$ $480\times 854$ $72.9$ $75.5$ $70.4$ $73.4$ $72.9$ $74.1$
${}^{{\dagger}}$ [9] semi-sup. $\checkmark$ $480\times 854$ 75.4 76.6 74.3 78.7 77.7 79.6
@@
Figure 4: Multi-object video segmentation on YTVOS18-m (left) and DAVIS17-m (right).
@@
${}^{\ddagger}$-flow [92] $\checkmark$ $128\times 224$ $72.1$ $67.6$ $70.0$
@@
${}^{\ddagger}$-TTA [92] $\checkmark$ $\checkmark$ $480\times 854$ $80.9$ $72.3$ $72.7$
@@
${}^{\ddagger}$urs $\checkmark$ $\checkmark$ $128\times 224$ 81.1 $76.6$ 81.9
@@
Table 2: Quantitative comparison on single object video segmentation benchmarks. “H.A.” indicates that human annotations are involved during training. normal-$\ddagger$ corresponds to the methods that rely on human-label-free supervision on synthetic data.
Figure 5: Single object segmentation on DAVIS16 (left), SegTrackv2 (middle), FBMS-59 (right).
4.3. Multiple object video segmentation
Our findings are further supported by comprehensive ablation studies provided in Appendix C. These studies verify our hyperparameter choices (e.g., the number of exemplar masks) and architectural designs (e.g., the choice of self-supervised features) during the synthetic training phase. Moreover, we detail how various self-supervised adaptation stages contribute to the overall performance.
Comparison to the Segment Anything Model (SAM) [38]. Recent studies highlight SAM as a robust general segmentor with diverse prompt options. To assess its efficacy for refining flow-based segmentation, we prompt SAM with masks predicted by the OCLR-flow model in a per-frame manner. As shown in Table 1, the resultant refinements (OCLR-flow + SAM) are not as effective as the refinements provided by our selection-correction model (Ours), owing to the limited utilisation of temporal consistency in the former approach. On the other hand, prompting SAM with masks predicted by our model does lead to improvements (Ours + SAM). This demonstrates that the SAM-based refinement is complementary to our method built on the persistence of appearance streams. Visualisations of the SAM-based refinements are provided in Figure 11 in the Appendix.
4.4. Single object video segmentation
@@ 5. Discussion
Nevertheless, some limitations have also been revealed under extremely complex motion scenarios: First, when a significant majority of problematic flow-predicted masks are present (e.g., due to a large proportion of stationary frames in the sequence), the selection of meaningful exemplar masks becomes challenging, consequently affecting the correction process. This problem could be potentially mitigated by incorporating additional appearance priors into the flow-based proposal stage. Second, the adopted frozen DINO ViT mainly encodes semantic-level information, which could result in difficulties in discerning entangled objects with similar appearances at the instance level. This highlights a possible research direction, aiming to develop instance-level appearance representations by leveraging flow-predicted information in complex multi-object scenes.
@@ Acknowledgements
Appendix A Architecture and implementation details
A.1 Appearance-based mask selector
@@
A.2 Appearance-based mask corrector
@@ A.3 Implementation details
Trainable parameters. The proposed appearance refinement models are lightweight, totaling $19.5M$ trainable parameters ($3.3M$ parameters for the mask selector and $16.2M$ parameters for the mask corrector). During the self-supervised adaptation, we fine-tune the OCLR-flow model, which comprises $20.7M$ trainable parameters.
@@
Inference time analysis. The overall inference time of the model is summarised in Table 14, where the inference time taken for a forward pass (including the flow-based proposal and the appearance-based refinement) is $0.098s$ per frame.
@@
Appendix B Datasets
B.1 Synthetic training dataset
@@ B.2 Evaluation datasets
YouTubeVOS2018 [93] (short form “YTVOS18”) is a large-scale multi-object video segmentation dataset consisting of $4453$ long-video sequences with around $200$k annotations sparsely provided at every $5^{\text{th}}$ frame. In particular, YTVOS18 is designed as a benchmark for semi-supervised VOS tasks (i.e., first-frame groundtruth provided). As a result, some sequences would be not suitable for motion segmentation evaluation due to problems such as common motion, unannotated moving objects, etc. Instead, we select a subset of YTVOS18 to form a new multi-object motion segmentation benchmark, named YouTubeVOS2018-motion (short form “YTVOS18-m”).
B.3 YouTubeVOS2018-motion formulation
The YTVOS18-m dataset is manually curated from the YTVOS18 training split and composed of $120$ selected sequences, with each sequence containing up to three moving objects. Compared to the DAVIS sequences, the selected YTVOS sequences comprise more realistic situations with variable motion patterns including a mixture of partial motion, background distraction, object occlusion and interactions, etc. We summarise our selection criteria into three major aspects regarding appearance, annotations and object motion, as outlined below.
@@ B.3 YouTubeVOS2018-motion formulation
Appendix C Ablation study
In this section, we present comprehensive ablation studies to validate our model performance on multi-object benchmarks. Specifically, in Sec. C.1, we focus on the synthetic pre-training stage, verifying hyperparameter choices and key design elements of our proposed models. Sec. C.2 provides detailed results at various stages of self-supervised adaptation. This includes self-supervised training of the mask corrector and the adaptation of the flow segmentation model. In Sec. C.3, we investigate how our self-supervised training scheme effectively mitigates the Sim2Real disparity introduced by the upsampling process.
C.1 Pre-training on synthetic videos
In Sec. C.1.1, we report the results of our synthetically-trained mask selector and compare it with other selection mechanisms. Sec. C.1.2 provides the comparison of various model configurations for the mask corrector during the synthetic pre-training stage.
C.1.1 Synthetically-trained mask selection
To evaluate the selection accuracy, we propose a metric that assesses the ability of the selector to identify $t=10$ high-quality exemplar masks per object while avoiding the problematic ones. Specifically, for each object, we rank the frames in the sequence based on the intersection-over-union (IoU) values between the flow-predicted masks and the groundtruth segmentation masks, from which the top $10$ frames and the bottom $10$ frames are extracted to form a set of $20$ frames. The selection accuracy is then measured by the Average Precision (AP) of selecting the top $10$ frames from this set.
@@ C.1.1 Synthetically-trained mask selection
C.1.2 Synthetically-trained mask correction
@@
$\checkmark$         $\textbf{60.9}{\scriptstyle\;{\color[rgb]{1.,.4,.4}+14.4}}$        $\textbf{60.0}{\scriptstyle\;{\color[rgb]{1.,.4,.4}+5.5}}$
@@
$\checkmark$ $0.098$        $64.6{\scriptstyle\;{\color[rgb]{1.,.4,.4}+3.7}}$        $63.2{\scriptstyle\;{\color[rgb]{1.,.4,.4}+3.2}}$
$\checkmark$ $\checkmark$ $0.098$        $66.6{\scriptstyle\;{\color[rgb]{1.,.4,.4}+2.0}}$        $65.4{\scriptstyle\;{\color[rgb]{1.,.4,.4}+2.2}}$
$\checkmark$ $\checkmark$ $\checkmark$ $2.030$        $66.9{\scriptstyle\;{\color[rgb]{1.,.4,.4}+0.3}}$        $66.3{\scriptstyle\;{\color[rgb]{1.,.4,.4}+0.9}}$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $5.600$        $\textbf{67.1}{\scriptstyle\;{\color[rgb]{1.,.4,.4}+0.2}}$        $\textbf{67.0}{\scriptstyle\;{\color[rgb]{1.,.4,.4}+0.7}}$
@@
C.1.3 Repeating (multiple run) experiments
C.2 Self-supervised adaptation to real-world videos
@@
C.3 Supervised adaptation: an investigation

### v2

@@ 3.4 Sim2Real Training Procedure
3.4.1 Pre-training on Synthetic Videos.
@@
$L_{\text{corrector}}=\lambda_{\text{mask}}L_{\text{mask}}+\lambda_{\text{feat-recon}}L_{\text{feat-recon}}$ (8)
@@
3.4.2 Self-Supervised Adaptation to Real-World Videos.
The Sim2Real domain gap is mitigated by self-supervised training on target real-world datasets. Specifically, for each real-world video dataset, we first employ our synthetically-trained selection-correction model to generate segmentation masks starting from the OCLR-flow proposals. The resultant masks are then treated as pseudo-labeling to fine-tune the mask corrector and the OCLR-flow segmentation model. We provide more details on each fine-tuning process below, and verify the effectiveness of each stage via ablation analysis in in Section 0.C.2.
Adapting the Mask Correction Model. The overall loss for fine-tuning remains the same as for synthetic training (Eq. 8). However, for adaptation, the refined masks generated by the synthetically-trained corrector (i.e., {~$\{\tilde{\mathcal{M}}_{i}^{k}\}$) are treated as pseudo-GT, and become the target masks in the loss. Note that, we only apply pseudo-labeling at the frames where exemplar masks present due to their high quality. Apart from the self-supervised training at the dataset level, we also employ a test-time per-sequence adaptation to further boost the performance. A detailed description is given in Section 0.A.3.
@@
4 Experiments
4.1 Datasets and Metrics
@@
Multi-Object Benchmark. We adopt the DAVIS2017-motion dataset (short form “DAVIS17-m”) that is proposed in [102] with jointly moving objects re-annotated. We further formulate a YouTubeVOS2018-motion dataset (short form “YTVOS18-m”), which is a subset of YTVOS18 [103], containing $120$ sequences with one or more dominant moving objects. Visualisations of multi-object sequences can be found in Figure 4. Additional information regarding this new benchmark is provided in Section 0.B.3.
@@
Evaluation Metrics. We benchmark our single object video segmentation performance using region similarity ($\mathcal{J}$). For multi-object video segmentation, we additionally report the contour similarity ($\mathcal{F}$[61].
4.2 Implementation Details
For pre-processing, we apply RAFT [87] to estimate optical flow with frame gaps ±$\pm 1$ (except for FBMS and YTVOS18-m with frame gaps ±$\pm 3$ owing to relatively slow video motions). All optical flows and RGB frame inputs are resized to $128\times 224$. The motion segmentation model proposes $K=3$ objects for each sequence, from which $t=10$ exemplar masks are selected for each object. Also, we use self-supervised pre-trained DINO transformers [11], ViT-S/16 for mask selector, ViT-S/8 for corrector, resulting in DINO feature resolutions $8\times 14$ and $16\times 28$, respectively. As a final post-processing step, we apply Conditional Random Field (CRF) segmentation to further refine the output masks. We refer the reader to Appendix 0.A for more details regarding the model architecture, training settings, and inference time analysis.
4.3 Multiple Object Video Segmentation
@@
VideoCutLER [97] un-sup. $\checkmark$ $480\times 854$ $57.0$ $59.0$ $55.1$ $57.3$ $57.4$ $57.2$
@@
CorrFlow [47] semi-sup. $\checkmark$ $480\times 854$ $55.2$ $60.0$ $50.4$ $54.0$ $54.2$ $53.7$
UVC [52] semi-sup. $\checkmark$ $480\times 854$ $66.9$ $70.4$ $63.4$ $65.5$ $66.2$ $64.7$
CRW [34] semi-sup. $\checkmark$ $480\times 854$ $72.9$ $75.5$ $70.4$ $73.4$ $72.9$ $74.1$
DINO [11] semi-sup. $\checkmark$ $480\times 854$ 75.4 76.6 74.3 78.7 77.7 79.6
@@
Figure 4: Qualitative comparison on multi-object video segmentation benchmarks, including YTVOS18-m (left) and DAVIS17-m (right).
Our findings are further supported by comprehensive ablation studies provided in Appendix 0.C. These studies verify our hyperparameter choices (e.g., the number of exemplar masks) and architectural designs (e.g., the choice of self-supervised features) during the synthetic training phase. Moreover, we detail how various self-supervised adaptation stages contribute to the overall performance.
Comparison to the Segment Anything Model (SAM) [43]. Recent studies highlight SAM as a robust general segmentor with diverse prompt options. To assess its efficacy for refining flow-based segmentation, we prompt SAM with masks predicted by the OCLR-flow model in a per-frame manner. As shown in Table 1, the resultant refinements (OCLR-flow + SAM) are not as effective as the refinements provided by our selection-correction model (Ours), owing to the limited utilisation of temporal consistency in the former approach. On the other hand, prompting SAM with masks predicted by our model does lead to improvements (Ours + SAM). This demonstrates that the SAM-based refinement is complementary to our method built on the persistence of appearance streams. Visualisations of SAM-based refinements are given in Appendix 0.E.
@@
OCLR-flow [102] $\checkmark$ $128\times 224$ $72.1$ $67.6$ $70.0$
DS [109] $\checkmark$ $\checkmark$ $240\times 426$ $79.1$ $72.1$ $71.8$
@@
OCLR-TTA [102] $\checkmark$ $\checkmark$ $480\times 854$ $80.9$ $72.3$ $72.7$
@@
Ours $\checkmark$ $\checkmark$ $128\times 224$ 81.1 $76.6$ 81.9
@@
Table 2: Quantitative comparison on single object video segmentation benchmarks. “H.A.” indicates that human annotations are involved during training. $\ddagger$ corresponds to methods that rely on human-label-free supervision on synthetic data.
4.4 Single Object Video Segmentation
@@
Figure 5: Qualitative comparison on single object video segmentation benchmarks, including DAVIS16 (left), SegTrackv2 (middle), and FBMS-59 (right).
@@ 5 Discussion
Nevertheless, some limitations have also been revealed under extremely complex motion scenarios: First, when a significant majority of problematic flow-predicted masks are present (e.g., due to a large proportion of stationary frames in the sequence), selecting meaningful exemplar masks becomes challenging, consequently affecting the correction process. This problem could be potentially mitigated by incorporating additional appearance priors into the flow-based proposal stage. Second, the adopted frozen DINO ViT mainly encodes semantic-level information, which could result in difficulties in discerning entangled objects with similar appearances at the instance level. This highlights a possible research direction, aiming to develop instance-level appearance representations by leveraging flow-predicted information in complex multi-object scenes.
@@ Acknowledgements
Appendix
This appendix contains the following parts:
  • Appendix 0.A: Architecture and Implementation Details, where we provide more information regarding architectural designs and implementation details.
  • Appendix 0.B: Datasets, where we provide a detailed description of all datasets adopted for training and evaluation.
  • Appendix 0.C: Ablation Studies, where we conduct a comprehensive ablation analysis of the key designs and components of the model.
  • Appendix 0.D: Additional Quantitative Results, where complete quantitative comparisons are provided between our method and prior works.
  • Appendix 0.E: Additional Qualitative Results are also provided on both final predictions and intermediate results.
  • Appendix 0.F: Potential Societal Impacts of our work are discussed.
Appendix 0.A Architecture and Implementation Details
0.A.1 Appearance-Based Mask Selector
@@
0.A.2 Appearance-Based Mask Corrector
@@ 0.A.3 Implementation Details
Trainable Parameters. The proposed appearance refinement models are light-weight, totaling $19.5M$ trainable parameters ($3.3M$ parameters for the mask selector and $16.2M$ parameters for the mask corrector). During the self-supervised adaptation, we fine-tune the OCLR-flow model, which comprises $20.7M$ trainable parameters.
@@
Inference Time Analysis. The overall inference time of the model is summarised in Table 14, where the inference time taken for a forward pass (including flow-based proposal and appearance-based refinement) is $0.098s$ per frame.
@@
Appendix 0.B Datasets
0.B.1 Synthetic Training Dataset
@@ 0.B.2 Evaluation Datasets
YouTubeVOS2018 [103] (short form “YTVOS18”) is a large-scale multi-object video segmentation dataset consisting of $4453$ long-video sequences with around $200$k annotations sparsely provided at every $5^{\text{th}}$ frame. In particular, YTVOS18 is designed as a benchmark for semi-supervised VOS tasks (i.e., first-frame groundtruth provided). As a result, some sequences would be not suitable for motion segmentation evaluation due to problems such as stationary objects, unannotated moving objects, etc. Instead, we select a subset of YTVOS18 to form a new multi-object motion segmentation benchmark, named YouTubeVOS2018-motion (short form “YTVOS18-m”).
0.B.3 YouTubeVOS2018-Motion Dataset
To curate a new evaluation benchmark for motion segmentation (where object discovery heavily relies on motion information), we manually select $120$ sequences, each containing up to three predominantly moving objects, from the YTVOS18 training split. Note, this dataset is only used for evaluation purposes. Compared to the DAVIS sequences, the selected YTVOS sequences comprise more realistic situations with variable motion patterns including a mixture of partial motion, background distraction, object occlusion and interactions, etc. We summarise our selection criteria into three major aspects regarding appearance, annotations and object motion, as outlined below.
@@ 0.B.3 YouTubeVOS2018-Motion Dataset
In summary, YTVOS18-m is introduced as a motion segmentation benchmark, featuring multi-object sequences characterized by dominant object movements. This stands in contrast to the original YTVOS18 dataset, as YTVOS18-m sequences are curated to support motion-based object discovery, offering a standardised benchmark for evaluating motion segmentation methods. Samples of YTVOS18-m sequences are given in Figures 12, 11 and 10.
Appendix 0.C Ablation Study
In this section, we present comprehensive ablation studies to validate our model performance on multi-object benchmarks. Specifically, in Sec. 0.C.1, we focus on the synthetic pre-training stage, verifying hyperparameter choices and key design elements of our proposed models. Sec. 0.C.2 provides detailed results at various stages of self-supervised adaptation. This includes self-supervised training of the mask corrector and the adaptation of the flow segmentation model. In Sec. 0.C.3, we investigate how our self-supervised training scheme effectively mitigates the Sim2Real disparity introduced by the upsampling process. Sec. 0.C.4 demonstrates that SAM-related models can be complementary with our model to achieve further performance improvement.
0.C.1 Pre-training on Synthetic Videos
In Sec. 0.C.1.1, we report the results of our synthetically-trained mask selector and compare it with other selection mechanisms. Sec. 0.C.1.2 provides the comparison of various model configurations for the mask corrector during the synthetic pre-training stage.
0.C.1.1 Synthetically-Trained Mask Selection.
To evaluate the selection accuracy, we propose a metric that assesses the ability of the selector to identify $t=10$ high-quality exemplar masks per object while avoiding the problematic ones. Specifically, for each object, we rank the frames in the sequence based on the IoU values between the flow-predicted masks and the groundtruth segmentation masks, from which the top $10$ frames and the bottom $10$ frames are extracted to form a set of $20$ frames. The selection accuracy is then measured by the Average Precision (AP) of selecting the top $10$ frames from this set.
@@ 0.C.1 Pre-training on Synthetic Videos
0.C.1.2 Synthetically-Trained Mask Correction.
@@
$\checkmark$         $\textbf{60.9}{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+14.4}}$        $\textbf{60.0}{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+5.5}}$
@@
0.C.2 Self-Supervised Adaptation to Real-World Videos
@@
$\checkmark$ $0.098$        $64.6{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+3.7}}$        $63.2{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+3.2}}$
$\checkmark$ $\checkmark$ $0.098$        $66.6{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+2.0}}$        $65.4{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+2.2}}$
$\checkmark$ $\checkmark$ $\checkmark$ $2.030$        $66.9{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+0.3}}$        $66.3{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+0.9}}$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $5.600$        $\textbf{67.1}{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+0.2}}$        $\textbf{67.0}{\scriptstyle\;{\color[rgb]{1.,.4,.4}\definecolor[named]{pgfstrokecolor}{rgb}{1.,.4,.4}+0.7}}$
@@
0.C.3 Supervised Adaptation: An Investigation